In 2019, X Company's profit function was 0.32R - $91,400, where R is revenue. In 2020, the relationship between revenue and variable costs will not change, but fixed costs will decrease by $21,936. Assuming a tax rate of 35%, what will revenue have to be in order for X Company to earn $45,600 after taxes in 2020?
Answer #1

Answer

>0.32 is the COntribution margin ratio, which is revenue - variable cost.

A After tax target profits $45,600
B = A/65% Before tax target profits $70,154
C = $ 91400 - 21936 Fixed Cost $69,464
D = B+C Total contribution margin required for target profits $139,618
E CM Ratio as per equation 0.32
F = D/E Revenue required to be earned for target profits $436,306 Answer